### v3

# Beyond Nonconvexity: A Universal Trust-Region Method with New Analyses

Abstract
The trust-region (TR) method is renowned historically for its robustness in nonconvex problems and extraordinary numerical performance, but the study of its performance in convex optimization is somehow limited. This paper complements the existing literature by presenting a universal trust-region method that simultaneously incorporates the quadratic regularization and ball constraint. In particular, we introduce a novel descent property tailored for trust-region-type algorithms, enabling us to unify and streamline the analysis for both convex and nonconvex optimization. Our method exhibits an iteration complexity of $\tilde{O}(\epsilon^{-3/2})$ to find an $\epsilon$-approximate second-order stationary point for nonconvex optimization. Meanwhile, the analysis reveals that the universal method attains an $O(\epsilon^{-1/2})$ complexity bound for convex optimization. Finally, we develop an adaptive universal method to address practical implementations. The numerical results show the effectiveness of our method in both nonconvex and convex problems.
1 Introduction
Second-order methods are renowned for their faster convergence compared to first-order methods, and the capability to find second-order stationary points, see Cartis et al. [1], Nocedal and Wright [2], Nesterov [3], Carmon et al. [4, 5]. Among these methods, the trust-region (TR) method [6] stands out as a representative approach due to its robustness in nonconvex problems and extraordinary numerical performance. Many linear or nonlinear programming solvers are using the trust-region method as an important building block, to name a few, Knitro [7], IPOPT [8], and PDFO [9], etc. Also, in the machine learning field, it inspires the well-known trust-region policy optimization [10, 11].
Recall that the TR method operates on the local quadratic approximation of objective function $f:\mathbb{R}^{n}\to\mathbb{R}$ within a TR ball constraint:
$\min_{d\in\mathbb{R}^{n}}$ $m_{k}(d):=\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}\nabla^{2}f(x_{k})d,$ (1.1)
$\operatorname{s.t.}$ $\|d\|\leq\Delta_{k}.$
Once the TR step $d_{k}$ is generated, the TR method decides whether to update or adjust the radius $\Delta_{k}$ based on the so-called ratio test:
$\rho_{k}~{}=\frac{f(x_{k}+d_{k})-f(x_{k})}{m_{k}(d_{k})-m_{k}(0)}.$ (1.2)
Despite its success in practice as mentioned above, the theoretical development of the TR method is somehow incomplete.
For instance, results in the early literature [12, 13, 6] only established (asymptotic) local convergence to $\epsilon$-approximate second-order stationary points ($\epsilon$-SOSP). These results also implied that TR has a nonasymptotic iteration complexity of $O(\epsilon^{-2})$ for $\epsilon$-SOSPs. However, such convergence rate is no better than first-order methods [3] for $\epsilon$-approximate first-order stationary points ($\epsilon$-FOSP), and an improved convergence rate is not aware for convex functions.
In contrast, as another representative second-order method, the Cubic Regularized (CR) Newton method has the updating rule
$d_{k}^{\text{CR}}=\arg\min_{d\in\mathbb{R}^{n}}\ m^{\text{CR}}_{k}(d):=\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}\nabla^{2}f(x_{k})d+\frac{\sigma_{k}}{3}\|d\|^{3},\sigma_{k}>0.$ (1.3)
This method was originally proposed by Griewank [14] and further analyzed by Nesterov and Polyak [15]. Later, Cartis et al. [16, 17] proposed the adaptive version of this method. It shows a convergence rate of $O(\epsilon^{-3/2})$ to find $\epsilon$-SOSPs for nonconvex functions and $O(\epsilon^{-1/2})$ for convex case simultaneously.
However, classical TR exhibits a clear gap to these convergence properties. This arises possibly due to two aspects: first, the classic TR method does not fully exploit the curvature information of Hessian; second, the ratio test (1.2) does not explicitly quantify the desired decrease for the TR steps [18]; both are crucial in modern complexity analysis.
Therefore, a plethora of TR variants (e.g., [18, 19, 20, 21]) have been proposed to improve iteration complexity over the years. The fixed-radius variant by Luenberger and Ye [21] achieves a complexity of $O(\epsilon^{-3/2})$ for finding $\epsilon$-SOSPs by controlling the stepsize proportionally to the tolerance ${\epsilon}^{1/2}$. However, this variant tends to be conservative for practical applications. By designing a contraction and expansion mechanism, Curtis et al. [18] introduced the first adaptive trust-region method (TRACE) matching the $\tilde{O}(\epsilon^{-3/2})$ complexity for finding $\epsilon$-SOSPs, where $\tilde{O}$ means there is a logarithmic term hidden in the complexity. Later, Curtis et al. [19] proposed another simplified variant of [18] while retaining the same dependency on $\epsilon$. A notable recent trust-region method introduced by Hamad and Hinder [20] achieves $O(\epsilon^{-3/2})$ complexity and improves the complexity’s dependency over Lipschitz constants by putting together the upper and lower bounds on the stepsizes. However, their algorithm only converges to an $\epsilon$-FOSP. Generally, these novel variants transcend the classic ones to bring a per-step sufficient decrease
$f(x_{k}+d_{k})-f(x_{k})\leq-\Omega(\epsilon^{3/2}).$ (1.4)
As a result, $O(\epsilon^{-3/2})$ rate of convergence can be achieved. However, for TR-type algorithms, such sufficient decrease (1.4) is highly nontrivial, due to its heavy dependence on the posterior dual multiplier of the subproblem (1.1), which is difficult to control in advance. As a price to overcome this challenge, these algorithms rely on complicated feedback to adjust the dual multiplier and may need additional assumptions. Moreover, these TR-type methods are initially designed for nonconvex optimization, and the associated convergence analysis in convex optimization is largely missing. However, even in many nonconvex problems, the objective function quickly turns weakly convex near some stationary points, which makes the study of TR-type methods in convex optimization equally important.
We summarize the above TR variants and other mainstream second-order methods in Table 1. As an observation, existing research fails to unleash the full potential of trust-region methods: $(a)$ While all aforementioned TR variants have improved complexity guarantee in the nonconvex case, to the best of our knowledge, neither of the nonasymptotic analyses extends to the convex case. Thus, whether a TR method has a global $O(\epsilon^{-1/2})$ convergence rate in convex optimization remains open. $(b)$ On the practical side, the existing adaptive trust-region methods tend to be complicated compared to other second-order methods [15, 22, 23], embedding different subroutines and nested loops. There is a clear mismatch between the theoretical guarantees and conciseness of implementation, thus calling for a simpler trust-region method with state-of-the-art complexity rates.
[b]
Table 1: Mainstream Second-order Methods. The notation means no such results exist in the corresponding paper, and $\tilde{O}$ hides the logarithmic terms. $\nabla f,\ \nabla^{2}f$ in the last column denote the paper assumes Lipschitz continuity of the gradient and Hessian of the objective separately.
Algorithm   Nonconvex   Convex   Local   Assumptions
Standard Trust-Region Method [2, 24] $O(\epsilon^{-2})$ Quadratic $\nabla^{2}f$
TRACE [18, 25] $\tilde{O}(\epsilon^{-3/2})$ Quadratic $\nabla^{2}f,\ \nabla f$
CAT [20] $O(\epsilon^{-3/2})$ Quadratic $\nabla^{2}f$
Fixed Radius TR [21] $O(\epsilon^{-3/2})$ Quadratic $\nabla^{2}f$
TR-NCG [19] $\tilde{O}(\epsilon^{-3/2})$ Quadratic $\nabla^{2}f$
Gradient-Regularized Newton Method [26, 22] $O(\epsilon^{-1/2})$ Superlinear $\nabla^{2}f$
SOAN2C [27] $\tilde{O}(\epsilon^{-3/2})$ Quadratic $\nabla f,\nabla^{2}f$
Damped Newton Method [28] $O(\epsilon^{-1/2})$ Quadratic $\ddagger$
Cubic Regularized Newton Method [15] $O(\epsilon^{-3/2})$ $O(\epsilon^{-1/2})$ Quadratic $\nabla^{2}f$
UTR-Simple $\tilde{O}(\epsilon^{-3/2})$ $O(\epsilon^{-1/2})$ Superlinear $\nabla^{2}f$
UTR-Adaptive $\tilde{O}(\epsilon^{-3/2})$ $O(\epsilon^{-1/2})$ Quadratic $\nabla^{2}f$
  • $\dagger$ The method [27, 19] does not provide local convergence analysis. We believe this should be true following standard analysis.
  • $\ddagger$ The method [28] assumes the objective function is strictly convex and uses a stronger version of the self-concordance [29].
Contributions.
Given this background, we introduce a simple yet potent TR variant: a universal trust-region (UTR) method based on the following subproblem:
$\min_{d\in\mathbb{R}^{n}}$ $\frac{1}{2}d^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\left\|\nabla f(x_{k})\right\|^{1/2}I\right)d+\nabla f(x_{k})^{T}d$ (1.5)
s.t. $\|d\|\leq\Delta_{k}:=r_{k}\|\nabla f(x_{k})\|^{1/2},$
where $\sigma_{k},r_{k}>0$ are hyperparameters. The feature of the subproblem above is incorporating gradient regularization in the local quadratic model and setting the trust-region radius proportional to the square root of the gradient norm. As a result, the $d_{k}$ generated by (1.5) exhibits a powerful feature referred to as Function-or-Stationarity-Decrease (FOSD, Property 2.2), i.e., after one successful update, either the function value decreases sufficiently or the gradient norm shrinks linearly. Different from (1.4), this desirable property enables a concise and clean algorithm which does not need sophisticated inner loop to control the posterior multiplier $\lambda_{k}$, and the associated complexity results have the state-of-the-art dependency over both the tolerance $\epsilon$ and the Lipschitz constant, requiring only the most standard assumptions.
More specifically, with explicit information on the Hessian Lipschitz constant, we propose a simple strategy to choose parameters $(\sigma_{k},r_{k})$ (Strategy 2.1), and UTR converges to an $\epsilon$-FOSP within $\tilde{O}(\epsilon^{-3/2})$ iterations. Surprisingly, we also find that UTR exhibits an improved $O(\epsilon^{-1/2})$ complexity when the objective function is convex. To the best of our knowledge, this is the first nonasymptotic analysis of TR-type methods in convex optimization.
Under the guidance of the simple strategy, we develop an adaptive approach (Strategy 4.1) that is more practical without relying on the Lipschitz constant. By fully exploiting the curvature information of the Hessian, this adaptive approach provides a convergence guarantee with an iteration complexity of $\tilde{O}(\epsilon^{-3/2})$ for finding $\epsilon$-SOSPs.
Finally, in numerical experiments, our method is compared against ARC [16, 17], the newly proposed TR variant [20], and the recent gradient regularized Newton method [22], demonstrating the effectiveness of UTR.
2 The Universal Trust-Region Method
In this paper, we consider the following unconstrained optimization problem
$\min_{x\in\mathbb{R}^{n}}\quad f(x),$ (2.1)
where $f:\mathbb{R}^{n}\to\mathbb{R}$ is twice differentiable and is bounded below, i.e., $f^{*}:=\inf_{x\in\mathbb{R}^{n}}f(x)>-\infty$. For nonconvex $f$, we aim to find $\epsilon$-approximate stationary points defined as follows:
Definition 2.1.
A point $x\in\mathbb{R}^{n}$ is called an $\epsilon$-approximate first-order stationary point ($\epsilon$-FOSP) of (2.1) if it satisfies
$\|\nabla f(x)\|\leq O(\epsilon)$ (2.2)
Moreover, we call $x$ an $\epsilon$-approximate second-order stationary point ($\epsilon$-SOSP) of (2.1) if it satisfies (2.2) and the following condition
$\lambda_{\min}(\nabla^{2}f(x))\geq\Omega(-\sqrt{\epsilon}).$ (2.3)
For convex $f$, we aim to find the approximate solution such that
$f(x)-f^{*}\leq O(\epsilon).$ (2.4)
Throughout the paper, we adopt the following standard assumption on the objective function $f$, commonly used in the complexity analysis of second-order methods.
Assumption 2.1.
The Hessian $\nabla^{2}f(x)$ of the objective function is Lipschitz continuous with constant $M>0$, i.e.,
$\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq M\|x-y\|\quad\forall x,y\in\mathbb{R}^{n}.$ (2.5)
As a consequence, Assumption 2.1 implies the following results.
Lemma 2.1 (Lemma 4.1.1, Nesterov [3]).
If $f:\mathbb{R}^{n}\mapsto\mathbb{R}$ satisfies Assumption 2.1, then for all $x,y\in\mathbb{R}^{n}$, we have
$\left\|\nabla f(y)-\nabla f(x)-\nabla^{2}f(x)(y-x)\right\|\leq\frac{M}{2}\|y-x\|^{2}$ (2.6a)
$\left|f(y)-f(x)-\nabla f(x)^{T}(y-x)-\frac{1}{2}(y-x)^{T}\nabla^{2}f(x)(y-x)\right|\leq\frac{M}{6}\|y-x\|^{3}$ (2.6b)
2.1 Overview of the Method
Now we introduce the universal trust-region method in Algorithm 1.
Data: Initial point $x_{0}\in\mathbb{R}^{n}$
1 for $k=0,1,\ldots,T$ do
2        Adjust $\left(\sigma_{k},r_{k}\right)$ by a proper strategy;
3        Solve the subproblem (1.5) and obtain the direction $d_{k}$;
4        if $d_{k}$ is good enough then
5               Update $x_{k+1}=x_{k}+d_{k}$;
6        else
7               Go to Line 3;
8
9        end if
10
11 end for
Algorithm 1 A Universal Trust-Region Method (UTR)
In particular, at each iteration $k$, we solve the subproblem (1.5). The mechanism of our trust-region method is straightforward, comprising only three major steps: setting the appropriate parameters $\sigma_{k}$ and $r_{k}$ by some strategy, solving the trust-region subproblem (1.5), and updating the iterate whenever $d_{k}$ is good enough.
The crux of our method lies in the selection of proper parameters $\sigma_{k}$ and $r_{k}$. This choice guides the model (1.5) to generate good steps that meet favorable descent properties to establish our iteration complexity bounds, i.e., the number of iterations $T$ required to find the desired solutions. We define the following properties which are tailored for the universal trust-region method.
Property 2.1 (Monotone-Decrease).
The step decreases the value of the objective function, that is for each iteration $k$,
$f(x_{k}+d_{k})-f(x_{k})\leq 0.$ (2.7)
Property 2.2 (Function-or-Stationarity-Decrease).
For some $10<\xi<$ 1, $\kappa>0$, the step $d_{k}$ either decreases the value of the objective function or decreases the gradient norm sufficiently, that is for each iteration $k$,
$f(x_{k}+d_{k})-f(x_{k})\leq-\frac{\kappa}{\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}\ \text{ or }\ \|\nabla f(x_{k}+d_{k})\|\leq\xi\|\nabla f(x_{k})\|.$ (2.8)
For Property 2.1, it serves as a safeguard to ensure that the algorithm never discards the progress made in the previous iterations. Property 2.2 is novel in the sense that we do not force every iteration to bring about a sufficient decrease in the function value but instead allow a linear contraction of the norm of the gradient as an alternative. Consequently, our analysis is largely simplified compared to other TR variants and uses only the most standard assumptions. Moreover, our analysis is almost unified for both convex and nonconvex optimization.
Unlike the classic ratio test, we design a new mechanism (Strategy 2.1) to ensure monotonicity: when we know the Lipschitz constant $M$. We show that the ratio test can be removed and every step is a good step. Moreover, even if we do not know the constant, we can perform a simple search procedure to ensure monotonicity.
We later prove the complexity results by establishing (2.7) and (2.8), and their modifications for convex functions (Property 3.1) as well as adaptiveness in choosing the parameters (Property 4.1). Moreover, we give general principles where parameter selection can be designed based on the information available at hand.
2.2 Basic Properties of the Method
We present some preliminary analysis of our method. Similar to the standard trust-region method, the optimality conditions of (1.5) are provided as follows.
Lemma 2.2.
The direction $d_{k}$ is the solution of (1.5) if and only if there exists a dual multiplier $\lambda_{k}\geq 0$ such that
$\|d_{k}\|\leq r_{k}\|\nabla f(x_{k})\|^{1/2}$ (2.9a)
$\lambda_{k}\left(\|d_{k}\|-r_{k}\|\nabla f(x_{k})\|^{1/2}\right)=0$ (2.9b)
$\left(\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)d_{k}=-\nabla f(x_{k})$ (2.9c)
$\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\succeq 0.$ (2.9d)
The results are directly obtained from Theorem 4.1 in Nocedal and Wright [2], and we omit the proof for succinctness. In the remaining part of this paper, we use $(d_{k},\lambda_{k})$ to denote the primal-dual solution pair of the subproblem at iteration $k$. Accounting for the optimality conditions (2.9a)-(2.9d), we could establish the following lemmas, which provide an estimation for the objective function and the gradient norm at the next iterate.
Lemma 2.3.
Suppose that Assumption 2.1 holds and $(d_{k},\lambda_{k})$ satisfies the optimal conditions (2.9a)-(2.9d), we have
$f(x_{k}+d_{k})\leq f(x_{k})-\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\|d_{k}\|^{3}.$ (2.10)
Additionally, if $\lambda_{k}\neq 0$, it follows that
$f(x_{k}+d_{k})\leq f(x_{k})-\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)r_{k}^{3}\|\nabla f(x_{k})\|^{3/2}.$ (2.11)
Proof.
By the $M$-Lipschitz continuous property of $\nabla^{2}f(x_{k})$ and Lemma 2.2, we conclude
$f(x_{k}+d_{k})-f(x_{k})$ $\leq\nabla f(x_{k})^{T}d_{k}+\frac{1}{2}d_{k}^{T}\nabla^{2}f(x_{k})d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
$=-\left(\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\right)\|d_{k}\|^{2}-\frac{1}{2}d_{k}^{T}\nabla^{2}f(x_{k})d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
$=-\frac{1}{2}\left(\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\right)\|d_{k}\|^{2}$
$\qquad-\frac{1}{2}d_{k}^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
$\leq-\frac{1}{2}\left(\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\right)\|d_{k}\|^{2}+\frac{M}{6}\|d_{k}\|^{3}$
$=-\frac{1}{2}\left(\lambda_{k}/\|\nabla f(x_{k})\|^{1/2}+\sigma_{k}\right)\|\nabla f(x_{k})\|^{1/2}\|d_{k}\|^{2}+\frac{M}{6}\|d_{k}\|^{3}$
$\leq-\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\|d_{k}\|^{3}.$
In the above, the first inequality comes from (2.6b); the first equality and the second inequality are due to the optimal properties (2.9c) and (2.9d), respectively; the last inequality is derived from (2.9a). As for the case $\lambda_{k}\neq 0$, the substitution $\|d_{k}\|=r_{k}\|\nabla f(x_{k})\|^{1/2}$ directly implies the validity of the inequality (2.11). ∎
Lemma 2.4.
Suppose that Assumption 2.1 holds and $(d_{k},\lambda_{k})$ satisfies the optimal conditions (2.9a)-(2.9d). If $\lambda_{k}=0$, then we have
$\|\nabla f(x_{k}+d_{k})\|\leq\left(\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}\right)\cdot\|\nabla f(x_{k})\|.$ (2.13)
Proof.
First, by the optimality condition (2.9c), when the dual variable $\lambda_{k}=0$, it follows that
$\|\nabla f(x_{k})+\nabla^{2}f(x_{k})d_{k}\|$ $=\left(\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\right)\cdot\|d_{k}\|$
$=\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\|d_{k}\|$
$=\sigma_{k}r_{k}\|\nabla f(x_{k})\|.$
With the Hessian Lipschitz continuity and Lemma 2.1, we get
$\|\nabla f(x_{k}+d_{k})\|$ $=\|\nabla f(x_{k}+d_{k})-\nabla f(x_{k})-\nabla^{2}f(x_{k})d_{k}+\nabla f(x_{k})+\nabla^{2}f(x_{k})d_{k}\|$
$\leq\|\nabla f(x_{k}+d_{k})-\nabla f(x_{k})-\nabla^{2}f(x_{k})d_{k}\|+\|\nabla f(x_{k})+\nabla^{2}f(x_{k})d_{k}\|$
$\leq\frac{M}{2}\|d_{k}\|^{2}+\sigma_{k}r_{k}\|\nabla f(x_{k})\|$
$\leq\frac{M}{2}r_{k}^{2}\|\nabla f(x_{k})\|+\sigma_{k}r_{k}\|\nabla f(x_{k})\|$
$\leq\left(\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}\right)\cdot\|\nabla f(x_{k})\|,$ (2.14)
where the second inequality is from (2.6a), and the third inequality is from (2.9a). ∎
Basic Principle of Choosing $(\sigma_{k},r_{k})$
The aforementioned Lemma 2.3 and Lemma 2.4 offer a valuable principle of selecting $\sigma_{k}$ and $r_{k}$ to guarantee that the step satisfies Property 2.1 and Property 2.2. In particular, it suffices to control
$~{}\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\cdot r_{k}^{3}>\frac{\kappa}{\sqrt{M}},\ \text{and}$ (2.15a)
$~{}\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}<\xi$ (2.15b)
for some $\kappa>0,\xi<1$. Thus, the choice of $\sigma_{k}$ and $r_{k}$ could be very flexible. For example, a vanilla approach can be constructed by discarding the term $\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}$ in (2.15a). Suppose the Lipschitz constant $M$ is given, we show that a strategy that fits (2.15) exists; namely, we can adopt a fixed rule of selecting $\sigma_{k}$ and $r_{k}$ as follows.
Strategy 2.1 (The Simple Strategy).
With the knowledge of Lipschitz constant $M$, we set
$\left(\sigma_{k},r_{k}\right)=\left(\frac{\sqrt{M}}{3},\frac{1}{3\sqrt{M}}\right)$ (2.16)
for all $k$ in Algorithm 1.
The universal trust-region method (Algorithm 1) equipped with such a simple choice reveals the following results.
Corollary 2.1.
Suppose Assumption 2.1 holds, by applying the Strategy 2.1, the steps generated by Algorithm 1 satisfy Property 2.1 and Property 2.2 with $\kappa=\frac{1}{81},\xi=\frac{1}{6}$, i.e.
$f(x_{k}+d_{k})\leq f(x_{k}).$
Furthermore, if the dual variable $\lambda_{k}\neq 0$, we have
$f(x_{k}+d_{k})-f(x_{k})\leq-\frac{1}{81\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}.$ (2.17)
While if the dual variable $\lambda_{k}=0$, we have
$\|\nabla f(x_{k}+d_{k})\|\leq\frac{1}{6}\|\nabla f(x_{k})\|.$ (2.18)
Proof.
Noticing $\lambda_{k}\geq 0$, it is easy to validate that
$\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\cdot r_{k}^{3}\geq\frac{1}{81\sqrt{M}}\quad\text{and}\quad\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}=\frac{1}{6}.$
Substituting the above inequalities into Lemma 2.3 and Lemma 2.4 completes the proof. ∎
In fact, one can adopt a choice of parameters without Lipschitz constants. Furthermore, if the information of $\lambda_{k}$ can be utilized, more aggressive strategies may be allowed. This direction is explored in the later sections to show stronger convergence to second-order stationarity. Nevertheless, the simple strategy (and a general design principle (2.15)) presented here is useful for understanding the building blocks of our method.
3 The Universal Trust-Region Method with a Simple Strategy
In this section, we give a convergence analysis of the universal method with the simple Strategy 2.1 to an $\epsilon$-approximate FOSP (see Definition 2.1) with an iteration complexity of $\tilde{O}\left(\epsilon^{-3/2}\right)$. The local convergence of this method is shown to be superlinear. Furthermore, the complexity can be further improved to $O\left(\epsilon^{-1/2}\right)$ for convex functions.
3.1 Global Convergence Rate for Nonconvex Optimization
For the nonconvex functions, we introduce the notation $x_{j_{f}}$ representing the first iterate satisfying
$\|\nabla f(x_{j_{f}})\|\leq\epsilon.$
We derive the convergence results according to Property 2.1 and Property 2.2. Based on the FOSD Property 2.2, we define the following index sets to classify the iteration generated by Algorithm 1,
$\mathcal{F}_{j_{f}}=\left\{k<j_{f}:f(x_{k}+d_{k})-f(x_{k})\leq-\frac{\kappa}{\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}\right\},\ \text{and}$ (3.1)
$\mathcal{G}_{j_{f}}=\left\{k<j_{f}:\|\nabla f(x_{k}+d_{k})\|\leq\xi\|\nabla f(x_{k})\|\right\}$
where $\kappa>0,~{}\xi<1$. From Corollary 2.1, we know each iteration in Algorithm 1 with Strategy 2.1 belongs to at least one of the above sets with $\kappa=\frac{1}{81}$ and $\xi=\frac{1}{6}$. If an iteration happens to belong to both, we assign it to set $\mathcal{F}_{j_{f}}$ for simplicity. Therefore, our goal is to provide an upper bound for the cardinality of sets $\mathcal{F}_{j_{f}}$ and $\mathcal{G}_{j_{f}}$. To begin with, we analyze $|\mathcal{F}_{j_{f}}|$ by evaluating the decrease in function value.
Lemma 3.1.
For Algorithm 1 with Strategy 2.1, suppose that Assumption 2.1 holds, then for any $k\in\mathcal{F}_{j_{f}}$, the function value decreases as
$f(x_{k+1})-f(x_{k})\leq-\frac{\kappa}{\sqrt{M}}\epsilon^{3/2}.$ (3.2)
Proof.
Note that for any $k\in\mathcal{F}_{j_{f}}$, the iterate $x_{k}$ satisfies $\|\nabla f(x_{k})\|>\epsilon,$ and hence this lemma is directly implied by the definition of $\mathcal{F}_{j_{f}}$. ∎
Based on Lemma 3.1, the upper bound regarding the cardinality of the set $\mathcal{F}_{j_{f}}$ is presented below.
Corollary 3.1.
Suppose that Assumption 2.1 holds, then the index set $\mathcal{F}_{j_{f}}$ satisfies
$|\mathcal{F}_{j_{f}}|\leq\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\epsilon^{-3/2}.$ (3.3)
Proof.
By Property 2.1, we know that Algorithm 1 is monotonically decreasing. By accumulating the function decrease (3.2), we have
$\frac{\kappa|\mathcal{F}_{j_{f}}|}{\sqrt{M}}\epsilon^{3/2}\leq\sum_{k\in\mathcal{F}_{j_{f}}}\frac{\kappa}{\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}\leq f(x_{0})-f^{*}.$
By rearranging items, we get the desired result. ∎
Now, it remains to establish an upper bound on the index set $|\mathcal{G}_{j_{f}}|$. We first show that the norms of the gradient of the iterates are uniformly bounded, which is one of the most important technical lemma for the analysis. It is highly nontrivial in TR-type methods due to the fact that the multiplier $\lambda_{k}$ is posterior and very hard to control in advance. We show that either it is bounded or it will bring a substantial amount of decrease in function value. As far as we know, this is a new result in the analyses for TR-type method. Since the proof is quite technical, we defer it to Appendix A.
Lemma 3.2.
Denote the sequence generated by Algorithm 1 with Strategy 2.1 as $\{x_{k}\}$, suppose that Assumption 2.1 holds, then for all $k\leq j_{f}$:
$\|\nabla f(x_{k})\|\leq\max\left\{\frac{1}{\xi}\left(\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\right)^{2/3},\|\nabla f(x_{0})\|,\frac{12\sqrt{M}}{\sqrt{\epsilon}}\left(f(x_{0})-f^{*}\right)\right\}:=G.$ (3.4)
As a result, the cardinality of the index set $\mathcal{G}_{j_{f}}$ could be bounded in terms of $|\mathcal{F}_{j_{f}}|$.
Lemma 3.3.
Suppose that Assumption 2.1 holds, then the index set $\mathcal{G}_{j_{f}}$ satisfies
$|\mathcal{G}_{j_{f}}|\leq\log(1/\xi)\log(G/\epsilon)|\mathcal{F}_{j_{f}}|,$ (3.5)
where $G$ is defined in Lemma 3.2.
Proof.
First, for arbitrary index $j\in\mathcal{F}_{j_{f}}$, we denote $n_{j}$ to be the maximum difference of indices between the two consecutive iterates in $\mathcal{F}_{j_{f}}$ after $x_{j}$. Then the iterates between $x_{j}$ and $x_{j+n_{j}}$ must be in $\mathcal{G}_{j_{f}}$. By the defintion of $\mathcal{G}_{j_{f}}$ and Lemma 3.2, we have that
$\epsilon<\|\nabla f(x_{j+n_{j}})\|\leq\|\nabla f(x_{j})\|\xi^{n_{j}}\leq\xi^{n_{j}}G.$
Therefore, the upper bound for $n_{j}$ could be evaluated as follow
$n_{j}<\log(1/\xi)\log(G/\epsilon).$
As a consequence, we return to $\mathcal{F}_{j_{f}}$ within $\lceil{n_{j}}\rceil$ iterations, and the inequality (3.5) follows. ∎
Now we are ready to prove the complexity result.
Theorem 3.1.
Suppose that Assumption 2.1 holds, the universal trust-region method (Algorithm 1) with Strategy 2.1 takes
$O\left(\sqrt{M}(f(x_{0})-f^{*})\epsilon^{-3/2}\log(G/\epsilon)\right)$
iterations to find an $\epsilon$-approximate first-order stationary point.
Proof.
We only need to find an upper bound for the summation
$|\mathcal{G}_{j_{f}}|+|\mathcal{F}_{j_{f}}|.$
By combining the results from Corollary 2.1, Corollary 3.1 and Lemma 3.3, we can obtain the desired result. ∎
We would like to echo again that the results obtained in this subsection rely on Property 2.1 and Property 2.2 rather than a specific strategy of choosing $(\sigma_{k},r_{k})$. Using Strategy 2.1 in the algorithm can be seen as a special choice.
3.2 Minimizing Convex Functions
In this subsection, we show that the universal trust-region method can achieve the state-of-the-art $O(\epsilon^{-1/2})$ iteration complexity making it comparable to other second-order methods when the objective function enjoys convexity, see Nesterov and Polyak [15], Doikov et al. [30], Doikov and Nesterov [26], Mishchenko [22]. Before delving into the analysis, we impose an additional property in this case.
Property 3.1.
The norm of the gradient at the next iterate is upper bounded as
$\|\nabla f(x_{k}+d_{k})\|\leq\left(1/\xi\right)\cdot\|\nabla f(x_{k})\|,$ (3.6)
where $10<\xi<$ 1 is the same as that defined in Property 2.2.
The above property is a safeguard for the iterates so that the gradient is bounded even in the case where $\lambda_{k}\neq 0$, cf. (2.8). We can again establish the validness of such property by, for example, Strategy 2.1.
Lemma 3.4.
Suppose that Assumption 2.1 holds. For the convex objective function $f$, by applying Algorithm 1 with Strategy 2.1, the step $d_{k}$ satisfies Property 3.1.
Proof.
If $\lambda_{k}=0$, the result is obvious from Corollary 2.1. When $\lambda_{k}\neq 0$, by a similar argument in the proof of Lemma 2.4, we have
$\|\nabla f(x_{k}+d_{k})\|$ $\leq\frac{M}{2}\|d_{k}\|^{2}+\|\nabla f(x_{k})+\nabla^{2}f(x_{k})d_{k}\|$
$\leq\frac{M}{2}r_{k}^{2}\|\nabla f(x_{k})\|+\|\nabla f(x_{k})+\nabla^{2}f(x_{k})d_{k}\|$ (3.7)
$=\frac{M}{2}r_{k}^{2}\|\nabla f(x_{k})\|+\left\|\left(\lambda_{k}I+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I\right)d_{k}\right\|$
$\leq\frac{1}{18}\|\nabla f(x_{k})\|+\left\|\left(\lambda_{k}I+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I\right)d_{k}\right\|$
$\leq\frac{19}{18}\|\nabla f(x_{k})\|,$ (3.8)
where the first equality is from applying triangle to (2.6a), the second inequality is from (2.9a), the equality is from (2.9c), the third inequality is due to $r_{k}=\frac{1}{3\sqrt{M}}$ in Strategy 2.1, and the last inequality comes from the following argument:
$\|d_{k}\|$ $=\left\|\left(\nabla^{2}f(x_{k})+\lambda_{k}I+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I\right)^{-1}\nabla f(x_{k})\right\|$
$\leq\left\|\left(\nabla^{2}f(x_{k})+\lambda_{k}I+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I\right)^{-1}\right\|\cdot\|\nabla f(x_{k})\|$
$\leq\frac{\|\nabla f(x_{k})\|}{\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}}.$
The last line of the above formula is because $f(x)$ is convex and $\nabla^{2}f(x_{k})$ is postive semidefinite. From Corollary 2.1 we know that $\xi=\frac{1}{6}$ in this case, hence we finish the proof. ∎
Similar to the previous discussion, we see that Property 3.1 can be met easily, e.g., by introducing an additional inequality to bound $M/2\cdot r_{k}^{2}$ from above (cf. (3.8)). Consequently, it is clear that a pair $(r_{k},\sigma_{k})$ satisfying Property 3.1 exists, in alignment with the principle (2.15) described in the previous subsection. In the following, we present the improved convergence results for convex functions. To establish the convergence result, we assume the sublevel set is bounded, which is widely used in the literature (e.g. Nesterov and Polyak [15], Doikov et al. [30]).
Assumption 3.1.
The diameter of the sublevel set $\mathcal{L}_{f}:=\left\{x:f(x)\leq f\left(x_{0}\right)\right\}$ is bounded by some constant $D>0$, which means that for any $x$ satisfying $f(x)\leq f\left(x_{0}\right)$ we have $\left\|x-x^{*}\right\|\leq D$.
For the convex optimization, we still adopt the notation $x_{j_{f}}$ representing the first iterate satisfying (2.4). Recalling the definition of the index sets in (3.1), we provide an upper bound for the cardinality of $\mathcal{F}_{j_{f}}$ in the following lemma.
Lemma 3.5.
Suppose that Assumption 2.1 and Assumption 3.1 hold and the objective function is convex, for Algorithm 1 with Strategy 2.1 , the index set $\mathcal{F}_{j_{f}}$ satisfies
$|\mathcal{F}_{j_{f}}|\leq\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}},$ (3.9)
where $\tau=\frac{\kappa}{\sqrt{M}}$, $\kappa$ is defined in Property 2.2.
Proof.
Using a similar argument as in Corollary 3.1, we denote the index set $\mathcal{F}_{j_{f}}$ in ascending order as $\{j_{f}(1),...,j_{f}(i),...\}$. Since $j_{f}(i)\in\mathcal{F}_{j_{f}}$ and the sequence is monotone, we have that
$f(x_{j_{f}(i+1)})-f(x_{j_{f}(i)})$ $=f(x_{j_{f}(i+1)})-f(x_{j_{f}(i)+1})+f(x_{j_{f}(i)+1})-f(x_{j_{f}(i)})$ (3.10)
$\leq f(x_{j_{f}(i)+1})-f(x_{j_{f}(i)})$
$\leq-\tau\left\|\nabla f(x_{j_{f}(i)})\right\|^{3/2}$
$\leq-\tau\left(\frac{f(x_{j_{f}(i)})-f^{*}}{D}\right)^{3/2},$
where the last inequality comes from the convexity of $f$ and
$f(x_{j_{f}(i)})-f^{*}\leq\nabla f(x_{j_{f}(i)})^{T}(x_{j_{f}(i)}-x^{*})\leq\left\|\nabla f(x_{j_{f}(i)})\right\|D.$
Denoting $\delta_{i}=f(x_{j_{f}(i)})-f^{*}$, we have
$\frac{1}{\sqrt{\delta_{i+1}}}-\frac{1}{\sqrt{\delta_{i}}}$ $=\frac{\sqrt{\delta_{i}}-\sqrt{\delta_{i+1}}}{\sqrt{\delta_{i}}\sqrt{\delta_{i+1}}}$
$=\frac{\delta_{i}-\delta_{i+1}}{\sqrt{\delta_{i}}\sqrt{\delta_{i+1}}(\sqrt{\delta_{i}}+\sqrt{\delta_{i+1}})}$
$\geq\frac{\tau}{D^{3/2}}\frac{\delta_{i}^{3/2}}{\sqrt{\delta_{i}}\sqrt{\delta_{i+1}}\left(\sqrt{\delta_{i}}+\sqrt{\delta_{i+1}}\right)}$
$\geq\frac{\tau}{2D^{3/2}},$
where the first inequality is due to (3.10) and the second inequality is from the monotonicity of the sequence. By summing over $i$ from $1$ to $k$, we obtain
$\frac{1}{\sqrt{\delta_{k}}}\geq\frac{1}{\sqrt{\delta_{k}}}-\frac{1}{\sqrt{\delta_{0}}}\geq\frac{k\tau}{2D^{3/2}},$
rearranging items yeilds
$\sqrt{\delta_{k}}\leq\frac{2D^{3/2}}{k\tau}.$
In other words, for any $k\in\mathcal{F}_{j_{f}}$, if $k\geq\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}$ then we have $\delta_{k}\leq\epsilon$ and the algorithm stops as $x_{k}$ is already an approximate solution. Thus, we conclude that the inequality (3.9) holds. ∎
Now we are ready to prove the complexity result of convex optimization.
Theorem 3.2.
Suppose that Assumption 2.1 and Assumption 3.1 hold, for the convex objective function, the universal trust-region method (Algorithm 1 with Strategy 2.1) takes
$O\left(\sqrt{M}D^{3/2}\epsilon^{-1/2}+\log\left(\|\nabla f(x_{0})\|/\epsilon\right)\right)$
iterations to find a point satisfying (2.4).
Proof.
Denote $T_{\epsilon}=2\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}+\log\frac{1}{\xi}\log\frac{\|\nabla f(x_{0})\|}{\epsilon}$, where $\tau$ is defined in Lemma 3.5, and thus it is sufficient to show that $j_{f}\leq T_{\epsilon}$.
On one hand, from Property 2.1 and Lemma 3.5, the number of iterations belonging to the set $\mathcal{F}_{j_{f}}$ would not exceed $\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}$, otherwise we have
$f\left(x_{T_{\epsilon}}\right)-f^{*}\leq\epsilon.$
On the other hand, by Property 2.2 and Property 3.1, we could deduce that after at most $T_{\epsilon}$ iterations, the gradient norm can be evaluated as follow
$\left\|\nabla f(x_{T_{\epsilon}})\right\|\leq\|\nabla f(x_{0})\|\left(\frac{1}{\xi}\right)^{\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}}\xi^{T_{\epsilon}-\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}}=\|\nabla f(x_{0})\|\left(\frac{1}{\xi}\right)^{\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}}\xi^{\sqrt{\frac{4D^{3}}{\epsilon\tau^{2}}}+\log\frac{1}{\xi}\log\frac{\|\nabla f(x_{0})\|}{\epsilon}}\leq\epsilon,$
which also demonstrates that
$f\left(x_{T_{\epsilon}}\right)-f^{*}\leq\nabla f(x_{T_{\epsilon}})^{T}(x_{T_{\epsilon}}-x^{*})\leq\left\|\nabla f(x_{T_{\epsilon}})\right\|\cdot D\leq O\left(\epsilon\right).$
As a result, $f(x_{T_{\epsilon}})-f^{*}\leq O(\epsilon)$ holds and we conclude $j_{f}\leq T_{\epsilon}$. Therefore, the convergence results for Strategy 2.1 is derived by Lemma 3.4 and Corollary 2.1. ∎
Notably, this complexity result is novel as trust-region methods have traditionally focused on nonconvex optimization problems, which closes the gap between the trust-region method and the cubic regularized Newton method for convex optimization.
3.3 Local Convergence
We now move to the local performance of Algorithm 1, we show that the method has superlinear local convergence when $\sigma_{k},r_{k}$ is updated as in Strategy 2.1. We first make a standard assumption in local analysis.
Assumption 3.2.
Denoting the sequence generated by the algorithm as $\{x_{k}\}$, we assume that $x_{k}\to x^{*}$, $k\to+\infty$, where $x^{*}$ satisfies
$\nabla f(x^{*})=0,\quad\nabla^{2}f(x^{*})\succeq\mu I\succ 0.$ (3.11)
First, we prove that under Assumption 3.2, when $k$ is sufficiently large, the trust-region constraint (2.9a) will be inactive in reminiscences of the classical results.
Lemma 3.6.
For Algorithm 1 with Strategy 2.1, suppose Assumption 3.2 holds, then the trust-region constraint (2.9a) will be inactive and $\lambda_{k}=0$ when $k\to+\infty$.
Proof.
Note that when $k\to+\infty$, by (2.9c) and (3.11), we have
$\|d_{k}\|$ $=\left\|\left(\nabla^{2}f(x_{k})+\frac{\sqrt{M}}{3}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)^{-1}\nabla f(x_{k})\right\|$ (3.12)
$\leq\frac{\|\nabla f(x_{k})\|}{\frac{\mu}{2}+\frac{\sqrt{M}}{3}\|\nabla f(x_{k})\|^{1/2}}$
$<r_{k}\|\nabla f(x_{k})\|^{1/2},$
where the last line is from $\mu\geq 6\sqrt{M}\|\nabla f(x_{k})\|^{1/2}$ when $k$ is large enough. This means $d_{k}$ is in the trust region, by (2.9b) we have $\lambda_{k}=0$, hence we finished the proof. ∎
A consequence of the above result is that the iterate gradually reduces to a gradient regularized Newton step for large enough $k$ in solving (1.5):
$d_{k}=-\left(\nabla^{2}f(x_{k})+\frac{\sqrt{M}}{3}\|\nabla f(x_{k})\|^{1/2}I\right)^{-1}\nabla f(x_{k}).$ (3.13)
Now we are ready to prove the local superlinear convergence of our algorithm.
Theorem 3.3.
Under Assumption 2.1 and Assumption 3.2, when $\sigma_{k},r_{k}$ are updated as in Strategy 2.1, Algorithm 1 has superlinear local convergence.
Proof.
Note that Algorithm 1 will recover the gradient regularized Newton method in the local phase since the problem becomes unconstrained, then it converges superlinearly, see Mishchenko [22]. ∎
4 The Adaptive Universal Trust-Region Method
In the above sections, we have provided a concise analysis of the universal trust-region method that applies uniformly to both convex and nonconvex optimization. Nevertheless, the limitation of Property 2.2 lies in its reliance on the unknown Lipschitz constant, which brings the challenge for efficient implementation. To enhance the capability of our method in practice, we provide an adaptive universal trust-region method (Algorithm 2), we show that it can preserve the global complexity of the simple strategy and find $\epsilon$-SOSPs. Also, the local convergence can be improved to quadratic.
4.1 The Adaptive Framework
To avoid using the Lipschitz constant $M$, several revisions should be made to our previous strategies of accepting the directions and tuning the parameters. In Algorithm 2, we impose an inner loop, indexed by $j$, for $(\sigma^{(j)}_{k},r^{(j)}_{k})$ parameterized by $\rho_{k}^{(j)}$. We terminate the loop $j$ until the iterates satisfy Property 4.1, which is defined as follows.
Data: Initial point $x_{0}\in\mathbb{R}^{n}$, tolerance $\epsilon>0$, decreasing constant $0<\eta<\frac{1}{32}$, $\frac{1}{4}<\xi<1$, initial penalty $\rho_{0}>0$, minimal penalty $\rho_{\min}>0$, penalty increasing parameter $\gamma_{1}>1$, penalty decreasing parameter $\gamma_{2}>1$;
1 for $k=0,1,\ldots,T_{\text{out}}$ do
2        Set $\rho_{k}^{(0)}=\rho_{k}$;
3        for $j=0,1,\ldots,T_{\text{in}}$ do
4               Update $\sigma_{k}^{(j)},r_{k}^{(j)}$ using Strategy 4.1;
5               Solve the trust-region subproblem (1.5) and obtain the direction $d_{k}^{(j)}$;
6               if Property 2.1 and Property 4.1 hold then
7                      Break;
8
9               else
10                      $\rho_{k}^{(j+1)}=\gamma_{1}\rho_{k}^{(j)}$;
11
12               end if
13
14        end for
15       Update $x_{k+1}=x_{k}+d_{k}^{(j)}$, $\rho_{k+1}=\max\{\rho_{\min},\rho_{k}^{(j)}/\gamma_{2}\}$;
16
17 end for
Algorithm 2 An Adaptive Universal Trust-Region Method
Property 4.1.
Given $\frac{1}{4}<\xi<1$, $0<\eta<\frac{1}{32}$, the step $d_{k}^{(j)}$ satisfies
{$f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}\text{ or }\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\epsilon^{3/2},\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon,\end{aligned}\right.$ (4.1)
where $\eta$ and $\rho_{k}^{(j)}$ are defined in Algorithm 2. If $f$ is nonconvex, the step $d_{k}^{(j)}$ additionally satisfies
$&\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\epsilon+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon.\end{aligned}\right.$ (4.2)
If $f$ is convex, the step $d_{k}^{(j)}$ should satisfy
$\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq 1/\xi\|\nabla f(x_{k})\|.$ (4.3)
Compared to Property 2.2, we allow no dependence on the Lipschitz constant $M$. The premise of this rule is that we can find a sufficiently large regularization $\sigma_{k}$ (or equivalently, sufficiently small $r_{k}$) based on Lemma 2.3 and Lemma 2.4 similar to other adaptive methods, see Cartis et al. [16], Curtis et al. [18], He et al. [31]. Besides, we proceed with the algorithm when the gradient norm is small so that one can find $\epsilon$-SOSPs.
As for the $(\sigma_{k}^{(j)},r_{k}^{(j)})$, we recall the principle (2.15a) that motivates the aforementioned simple strategy. As we directly ignore the curvature information, it only converges to a first-order stationary point when $f(x)$ is nonconvex. Therefore we propose the following adaptive strategy (Strategy 4.1) to allow convergence to second-order stationary points.
Strategy 4.1 (The Strategy for Second-order Stationary Points).
In the Line 5 of Algorithm 2, we apply the following strategy in Table 2.
Table 2: The Adaptive Strategy
Gradient Conditions Selection of $(\sigma_{k},r_{k})$
$\|\nabla f(x_{k})\|\geq\epsilon$ $\lambda_{\min}(\nabla^{2}f(x_{k}))\leq-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$ $\sigma_{k}^{(j)}=0,\ r_{k}^{(j)}=1/2\rho_{k}^{(j)}$
$\lambda_{\min}(\nabla^{2}f(x_{k}))\geq\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$
$-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$ $\sigma_{k}^{(j)}=\rho_{k}^{(j)},\ r_{k}^{(j)}=1/4\rho_{k}^{(j)}$
$\|\nabla f(x_{k})\|<\epsilon$ $\lambda_{\min}(\nabla^{2}f(x_{k}))>-\rho_{k}^{(j)}\epsilon^{1/2}$
$\lambda_{\min}(\nabla^{2}f(x_{k}))\leq-\rho_{k}^{(j)}\epsilon^{1/2}$ $\sigma_{k}^{(j)}=0,\ r_{k}^{(j)}=\epsilon^{1/2}/\left(2\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}\right)$
The symbol ✓means $x_{k}$ is already an $\epsilon$-SOSP and we can terminate the Algorithm 2.
We later justify that the direction $d_{k}^{(j)}$ will gradually be accepted at some $j$ (see Lemma 4.1). As shown in the next subsection, the adaptive method converges to $\epsilon$-SOSPs with the same complexity as the previous conceptual version. Furthermore, the adaptive version also allows us to adjust the regularization $\sigma_{k}$, which leads to a faster speed of local convergence. Certainly, such a strategy relies on additional information from the smallest eigenvalue. As the trust-region method very often utilizes a Lanczo-type method to solve the subproblems, e.g., Cartis et al. [1], Conn et al. [6], using the smallest eigenvalue of the Hessian incurs no significant cost, see Gratton et al. [27], Hamad and Hinder [20]. If instead we use a factorization-based method, the Cholesky factorization can also serve the purpose of the eigenvalue test: we may increase the dual-variable $\lambda_{k}$ if the factorization fails, in which case, an estimate of the smallest eigenvalue can be built from $\lambda_{k}$ and $\sigma_{k}$.
4.2 Global Convergence
In this subsection, we begin with the complexity analysis in the nonconvex case. We demonstrate that Algorithm 2 requires no more than $\tilde{O}\left(\epsilon^{-3/2}\right)$ iterations to converge to an $\epsilon$-approximate second-order stationary point satisfying (2.2) and (2.3). The following lemma shows that there exists an upper bound on the penalty parameter $\rho_{k}^{(j)}$, leading to the termination of the inner loop.
Lemma 4.1.
In Algorithm 2 with Strategy 4.1, there exists a uniform upper bound for the parameter $\rho_{k}^{(j)}$, that is
$\rho_{k}^{(j)}\leq\rho_{\max}:=\gamma_{1}\cdot\max\left\{\sqrt{\frac{M}{12(1-32\eta)}},\sqrt{\frac{M}{6(1-8\eta)}},\sqrt{\frac{M}{32\xi-8}},\sqrt{\frac{M}{8\xi}},\sqrt{\frac{M\xi}{8(1-\xi)}}\right\}.$ (4.4)
Since this lemma is quite technical, we delay the analysis in Appendix A. As a direct consequence of Lemma 4.1, the iteration number of the inner loop in Algorithm 2 could be upper bounded.
Corollary 4.1.
The number of oracle calls in inner $j$-loop of Algorithm 2 is bounded by $\log_{\gamma_{1}}\frac{\rho_{\max}}{\rho_{\min}}$.
Nonconvex Functions
Now we are ready to give a formal iteration complexity analysis of Algorithm 2. We show that for the nonconvex objective function with Lipschitz continuous Hessian, Algorithm 2 takes $\tilde{O}\left(\epsilon^{-3/2}\right)$ to find an $\epsilon$-approximate second-order stationary point $x$ satisfying (2.2) and (2.3).
Similar to the previous section, the following analysis is standard. As shown in Lemma 4.1, the iterates produced by Algorithm 2 will end up satisfying Property 4.1, therefore we can classify them into the following two categories
$\mathcal{F}_{p_{s}}$ =$=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\max\left\{\|\nabla f(x_{k})\|^{3/2},\epsilon^{3/2}\right\}\right\},$ (4.5)
$\mathcal{G}_{p_{s}}$ =$=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\},$
and denote $x_{p_{s}}$ as the first iteration satisfying (2.2) and (2.3). Then by the mechanism of the Algorithm 2, all indices belong to one of the sets defined in (4.5), and thus we only need to provide an upper bound for the summation
$T_{p_{s}}:=|\mathcal{F}_{p_{s}}|+|\mathcal{G}_{p_{s}}|.$ (4.6)
Similar to the previous section, we first show that the norm of the gradient of the iterates have an uniform upper bound, with the proof deferred to Appendix A.
Lemma 4.2.
Suppose Assumption 2.1 holds, denote the sequence generated by Algorithm 2 with Strategy 4.1 as $\{x_{k}\}$, then for all $k\leq p_{s}$:
f(xk)Ga,where
[…]
$G_{a}:=\max\left\{\left(\xi+\frac{M}{6\rho_{\min}^{2}}+16\eta\right)\left(\frac{\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\eta}\right)^{2/3},\|\nabla f(x_{0})\|,\;\frac{32\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\sqrt{\epsilon}}\right\}.$ (4.7)
For the index set $\mathcal{F}_{p_{s}}$ and $\mathcal{G}_{p_{s}}$, we conclude the following results.
Lemma 4.3.
Suppose that Assumption 2.1 holds, the cardinality of the index sets $\mathcal{F}_{p_{s}}$ and $\mathcal{G}_{p_{s}}$ satisfies
$|\mathcal{F}_{p_{s}}|\leq\frac{\rho_{\max}}{\eta}\left(f(x_{0})-f^{*}\right)\epsilon^{-3/2},$ (4.8)
and
$|\mathcal{G}_{p_{s}}|\leq\log(1/\xi)\log(G_{a}/\epsilon)|\mathcal{F}_{p_{s}}|.$ (4.9)
We omit the proofs as they are almost the same as Corollary 3.1 and Lemma 3.3. Therefore, we are ready to present the formal complexity result of Algorithm 2.
Theorem 4.1.
Suppose that Assumption 2.1 holds, Algorithm 2 takes
$O\left(\rho_{\max}(f(x_{0})-f^{*})\epsilon^{-3/2}\log\left(G_{a}/\epsilon\right)\log_{\gamma_{1}}\left(\rho_{\max}/\rho_{\min}\right)\right)$ (4.10)
iterations to find an $\epsilon$-approximate second-order solution satisfying (2.2) and (2.3).
Proof.
The result is directly implied by Lemma 4.3 and Corollary 4.1. ∎
Convex Functions
For the case where the objective function is convex, we are satisfied with points such that the norm of the gradient less than $\epsilon$, then the two sets becomes
$\mathcal{F}_{p_{s}}$ =$=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}\right\},$
$\mathcal{G}_{p_{s}}$ =$=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\}.$
The following analysis is the same as Algorithm 1 with Strategy 2.1. We directly provide the convergence result here.
Theorem 4.2.
Suppose that $f(x)$ is convex, Assumption 2.1 and Assumption 3.1 hold, then Algorithm 2 takes
$O\left(\rho_{\max}(f(x_{0})-f^{*})\log_{\gamma_{1}}\left(\rho_{\max}/\rho_{\min}\right)\epsilon^{-1/2}\right)$ (4.11)
iterations to find an $\epsilon$-approximate solution satisfying (2.4).
4.3 Local Convergence
In this subsection, we give the local performance of Algorithm 2 with Strategy 4.1 under Assumption 3.2, and show that the method has a local quadratic rate of convergence when $(\sigma_{k},r_{k})$ is updated as in Strategy 4.1.
Since in the local scenario, we can regard $\epsilon\to 0$, and $\rho_{k}^{(j)}$ has a uniform upper bound, then Strategy 4.1 will persist in the case when $k$ is sufficiently large:
$\lambda_{\min}(\nabla^{2}f(x_{k}))\geq\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2},$
in which we always set $(\sigma_{k}^{(j)},r_{k}^{(j)})=(0,1/2\rho_{k}^{(j)})$. The rest of the cases are irrelevant to our discussion. Similar to the previous discussion, we show that when $k$ is sufficiently large, the trust-region constraint (2.9a) will be inactive.
Lemma 4.4.
Suppose Assumption 2.1 and Assumption 3.2 hold, then the trust-region constraint (2.9a) will be inactive and $\lambda_{k}=0$ when $k\to+\infty$.
Proof.
Note that by (2.9c) and (3.11), we have
$\|d_{k}\|=\left\|\left(\nabla^{2}f(x_{k})+\lambda_{k}I\right)^{-1}\nabla f(x_{k})\right\|\leq\frac{\|\nabla f(x_{k})\|}{\|\nabla^{2}f(x_{k})+\lambda_{k}I\|}\leq\frac{2\|\nabla f(x_{k})\|}{\mu}$ (4.12)
when $k$ is sufficiently large. Also by (3.11), we know that there exist a constant $k_{l}>0$ bounded from above, such that for all $k\geq k_{l}$. Thus, we have
$\|\nabla f(x_{k})\|<\frac{\mu^{2}}{16\rho_{\max}^{2}},$
and then
$\|d_{k}\|\leq\frac{2\|\nabla f(x_{k})\|}{\mu}<\frac{\|\nabla f(x_{k})\|^{1/2}}{2\rho_{\max}}\leq r_{k}\|\nabla f(x_{k})\|^{1/2}.$
This means $d_{k}$ is in the trust region, by (2.9b) we have $\lambda_{k}=0$, and this completes the proof. ∎
As we set $\sigma_{k}=0$ when $k$ is sufficiently large, the step that solves (1.5) is equivalent to a Newton step $d_{k}=-\nabla^{2}f(x_{k})^{-1}\nabla f(x_{k})$ rather than a regularized Newton step, indicating the local quadratic convergence of our algorithm.
Theorem 4.3.
Suppose Assumption 2.1 and Assumption 3.2 hold, Algorithm 2 with Strategy 4.1 has quadratic local convergence.
Proof.
Note that as in the previous section, Algorithm 2 will recover the Newton method in the local phase, then it converges quadratically, see Theorem 3.5, Nocedal and Wright [2]. ∎
5 Numerical Experiments
In this section, we present numerical experiments. We implement the adaptive UTR (Algorithm 2) in Julia programming language. All experiments are performed on a MacBook Pro with Apple M2 Chip and 24GB LPDDR5 Memory. To enable efficient routines for trust-region subproblems, we implement two options. The first option utilizes the standard Cholesky factorization [2, Algorithm 4.3] and uses a hybrid bisection and Newton method to find the dual variable [32, 33]. When using this option, we name the method after UTR. The second option is an indirect method (so it is referred to as iUTR) by Krylov subspace iterations, which is consistent with the open-source implementation [34] of classical trust-region method and adaptive cubic regularized Newton method. Motivated from Curtis and Wang [25] and Cartis et al. [1, Chapter 10], we use the Lanczos method with inexactness of subproblem solutions.
CUTEst benchmarks
We conduct experiments on unconstrained problems with dimension $n\leq 5000$ in the CUTEst benchmarks [35]. Since many of these problems are nonconvex, we focus on comparisons with the classical trust-region method [6] and adaptive cubic regularized Newton method [16]. All methods use Krylov approaches to solve subproblems. Specifically, the classical trust-region method uses the Steihaug-Toint conjugate gradient method. Since both the classical trust-region method (Newton-TR-STCG) and adaptive cubic regularized method (ARC) are well studied, we directly use the implementation in Dussault [36].
We present our results in Table 3. We report $\overline{t}_{G},\overline{k}_{G}$ as scaled geometric means of running time in seconds and iterations (scaled by 1 second and 50 iterations, respectively). We regard a successful instance if it is solved within 200 seconds with an iterate $x_{k}$ such that $\|\nabla f(x_{k})\|\leq 10^{-5}$. If an instance fails, its iteration number and running time are set to $20,000$. We set the total number of successful instances as $\mathcal{K}$. Then we present the number of function evaluations and gradient evaluations by $\overline{k}_{G}^{f}$ and $\overline{k}_{G}^{g}$, respectively, where $\overline{k}_{G}^{g}$ also includes the Hessian-vector evaluations.
Table 3: Performance of different algorithms on the CUTEst dataset. $\overline{t}_{G},\overline{k}_{G},\overline{k}_{G}^{f},\overline{k}_{G}^{g}$ are computed as geometric means.
method $\mathcal{K}$ $\overline{t}_{G}$ $\overline{k}_{G}$ $\overline{k}_{G}^{f}$ $\overline{k}_{G}^{g}$
ARC 167.00 5.32 185.03 185.03 888.35
Newton-TR-STCG 165.00 6.14 170.44 170.44 639.64
iUTR 181.00 4.23 90.00 107.19 1195.47
In Table 3, iUTR has the most successful numbers, best running time as well as iteration performance. These results match the complexity analysis that unveils the benefits of gradient norm in both trust-region radii and regularization terms.
Logistic regression
For convex optimization, we test on logistic regression with $\ell_{2}$ penalty,
$f(x)=\frac{1}{N}\sum_{i=1}^{N}\log\left(1+e^{-b_{i}\cdot a_{i}^{T}x}\right)+\frac{\gamma}{2}\|x\|^{2}.$ (5.1)
where $a_{i}\in\mathbb{R}^{n},~{}b_{i}\in\{-1,1\}$, $i=1,2,\cdots,N.$ We set $\gamma=10^{-8}$ so that the Newton steps may fail at degenerate Hessians. Since the problem is convex, we focus on comparisons with the adaptive Newton method with cubics (ArC, Cartis et al. [16]) and variants of the regularized Newton method [22]. We implement the regularized Newton method (RegNewton) with fixed regularization $\sigma_{k}\in\{10^{-3},5\times 10^{-4}\}$ and an adaptive version following Mishchenko [22, Algorithm 2.3] that is named after RegNewton-AdaN+.
Figure 1: Logistic regression on LIVSVM instances: a4a (left) and w8a (right).
In Figure 1, we profile the performance of these methods in minimizing the gradient norm. The results show that the adaptive universal trust-region method is comparable to ArC and RegNewton-AdaN+, implying its competence in minimizing the convex functions.
Matrix Completion
A final experiment was conducted for matrix completion problems as a comparison to a recent TR method, CAT [20]. Briefly, the matrix completion task involves reconstructing a partially observed matrix $D\in\mathbb{R}^{n_{1}\times n_{2}}$. The goal is to recover the original dataset $D$ by a partially observed set $\Omega=\{(i,j)\}$ and the low-rank representation $D=PQ^{T}$, where $P\in\mathbb{R}^{n_{1}\times r}$ and $Q\in\mathbb{R}^{n_{2}\times r}$ with $r\ll n_{1}$ and $r\ll n_{2}$. When predicting the power consumption, $n_{1},n_{2}$ are number of features and data points, respectively. Allow a baseline regression model:
$d_{i,j}=\mu+r_{i}+c_{j},$ (5.2)
where $\mu$ represents the average of all observed measurements, $r_{i},c_{j}$ captures the deviations from $i$ and $j$. We focus on the following regularized problem:
$\min_{r,c,p,q}\sum_{(i,j)\in\Omega}(D_{i,j}-\mu-r_{i}-c_{j}-p_{i}q_{j}^{T})^{2}+\lambda(r_{i}^{2}+c_{j}^{2})+\lambda(\|p_{i}\|^{2}_{2}+\|q_{j}\|^{2}_{2})$
to fit the observed data under (5.2). The dataset used for this study is sourced from Ausgrid [20] that deals with the matrix size of $48\times 30$, $r=9$, all algorithms are terminated at $\|\nabla f(x_{k})\|\leq 10^{-7}$. We vary $\lambda\in[10^{-2},10^{-3},10^{-4}]$ to test the robustness under small regularizations. The results are presented in Table 4. While smaller $\lambda$ induces harder instances, iUTR has the best performance among competing algorithms in all instances. Unfortunately, the CAT method only converges to inaccuracy solution with $\|\nabla f(x_{k})\|\leq 10^{-5}$; hence we only include its results when $\lambda=10^{-2}$.
Table 4: Performance of different algorithms on Matrix Completion [20]. Here, we present the average iteration number, function, and gradient evaluations of five runs. We limit the run time to 1,000 seconds.
$\lambda$ method $\bar{k}$ $\bar{k}^{f}$ $\bar{k}^{g}$
$10^{-2}$ CAT 212 213 213
ARC 81 69 52
Newton-TR-STCG 80 81 55
iUTR 39 51 39
$10^{-3}$ ARC 271 152 121
Newton-TR-STCG 271 271 171
iUTR 78 91 78
$10^{-4}$ ARC 498 501 315
Newton-TR-STCG 502 316 238
iUTR 164 187 164
Appendix A Technical Proofs
Proof to Lemma 3.2
Proof.
Our proof is based on mathematical induction. Note that for all $k<j_{f}$, $\|\nabla f(x_{k})\|>\epsilon$, (3.4) holds for $i=0$. Suppose that (3.4) also holds for $i=k<j_{f}$, we will show that it holds for $i=k+1\leq j_{f}$. Since we use Strategy 2.1 and (2.15b) holds, from the second-order Lipschitz continuity of the Hessian matrix, we have
$\|\nabla f(x_{k+1})\|$ $\leq\frac{M}{2}\|d_{k}\|^{2}+\|\nabla f(x_{k})+\nabla^{2}f(x_{k})d_{k}\|$ (A.1)
$\leq\frac{M}{2}r_{k}^{2}\|\nabla f(x_{k})\|+\|\left(\lambda_{k}I+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I\right)d_{k}\|$
$\leq\left(\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}\right)\|\nabla f(x_{k})\|+\lambda_{k}\|d_{k}\|$
$\leq\xi\|\nabla f(x_{k})\|+\frac{1}{3\sqrt{M}}\lambda_{k}\|\nabla f(x_{k})\|^{1/2}.$
Let us discuss according to the value of $\|\nabla f(x_{k})\|$. First, suppose that $\|\nabla f(x_{k})\|\geq\left(\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\right)^{2/3}$, then we must have $\lambda_{k}=0$ and
$\|\nabla f(x_{k}+d_{k})\|\leq\xi\|\nabla f(x_{k})\|.$
Otherwise, we have $\lambda_{k}>0$ and from (3.2) it holds that
$f(x_{k}+d_{k})\leq f(x_{k})-\frac{\kappa}{\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}\leq f(x_{0})-\frac{\kappa}{\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}<f^{*}.$
We come to a contradiction and thus the claim holds.
Now consider the case $\epsilon\leq\|\nabla f(x_{k})\|<\left(\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\right)^{2/3}$. In this case if $\lambda_{k}=0$, (3.4) obviously holds, then we go to the case $\lambda_{k}>0$. We echo that (2.11) and Strategy 2.1 states that
$f(x_{k}+d_{k})$ $\leq f(x_{k})-\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)r_{k}^{3}\|\nabla f(x_{k})\|^{3/2}$
$\leq f(x_{k})-\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}\right)r_{k}^{3}\|\nabla f(x_{k})\|^{3/2}$
$\leq f(x_{k})-\frac{1}{18M}\lambda_{k}\|\nabla f(x_{k})\|.$
Therefore we have
$\lambda_{k}\|\nabla f(x_{k})\|\leq 18M\left(f(x_{0})-f^{*}\right).$ (A.2)
Plugging (A.2) into (A.1), we have
$\|\nabla f(x_{k}+d_{k})\|\leq\xi\|\nabla f(x_{k})\|+6\sqrt{M}\|\nabla f(x_{k})\|^{-1/2}\left(f(x_{0})-f^{*}\right).$ (A.3)
Consider function
$g(t)=\xi t+\frac{6\sqrt{M}}{\sqrt{t}}\left(f(x_{0})-f^{*}\right),t\in\left[\epsilon,\left(\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\right)^{2/3}\right).$
By taking derivative, we know that $g$ achieves its maximum at the bracket points of the interval, and consequently
$g(t)\leq\max\left\{\xi\epsilon+\frac{6\sqrt{M}}{\sqrt{\epsilon}}\left(f(x_{0})-f^{*}\right),\left(\xi+6\kappa^{1/3}\right)\left(\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\right)^{2/3}\right\}.$
Noting that from Strategy 2.1, $\xi=\frac{1}{6},\kappa=\frac{1}{81}$ and the fact that $\epsilon$ is a small quantity, we have $\xi\epsilon\leq\frac{6\sqrt{M}}{\sqrt{\epsilon}}\left(f(x_{0})-f^{*}\right)$, $\xi+6\kappa^{1/3}\leq\frac{1}{\xi}$, and thus
$g(t)\leq\max\left\{\frac{12\sqrt{M}}{\sqrt{\epsilon}}\left(f(x_{0})-f^{*}\right),\frac{1}{\xi}\left(\frac{\sqrt{M}}{\kappa}\left(f(x_{0})-f^{*}\right)\right)^{2/3}\right\}.$
Therefore from (A.3) we know that the claim holds for $i=k+1$ and the proof is finished. ∎
Proof to Lemma 4.1
Proof.
It is sufficient to show that for every $k$-th outer iteration, whenever the parameter $\rho_{k}^{(j)}$ satisfies
$\rho_{k}^{(j)}\geq\max\left\{\sqrt{\frac{M}{12(1-32\eta)}},\sqrt{\frac{M}{6(1-8\eta)}},\sqrt{\frac{M}{32\xi-8}},\sqrt{\frac{M}{8\xi}},\sqrt{\frac{M\xi}{8(1-\xi)}}\right\},$ (A.4)
the inner loop will terminate. Firstly, we consider the case where $\|\nabla f(x_{k})\|\leq\epsilon$ and $\lambda_{\min}(\nabla^{2}f(x_{k}))\leq-\rho_{k}^{(j)}\epsilon^{1/2}$. To facilitate the analysis, we introduce the concept of eigenpoint within the trust region:
$d_{k}^{E}:=\frac{\epsilon^{1/2}}{2\rho_{k}^{(j)}}v_{k}\;\;\mbox{such that}\;\;v_{k}^{T}\nabla f(x_{k})\leq 0,$ (A.5)
where $v_{k}$ is the unit eigenvector corresponding to the smallest eigenvalue $\lambda_{\min}(\nabla^{2}f(x_{k}))$. Note that for the eigenpoint $d_{k}^{E}$, it follows
$\nabla f(x_{k})^{T}d_{k}^{E}+\frac{1}{2}\left(d_{k}^{E}\right)^{T}\nabla^{2}f(x_{k})d_{k}^{E}\leq\frac{1}{2}\left(d_{k}^{E}\right)^{T}\nabla^{2}f(x_{k})d_{k}^{E}\leq-\frac{1}{8\rho_{k}^{(j)}}\epsilon^{3/2}.$
Moreover, since the eigenpoint is feasible, once the parameter $\rho_{k}^{(j)}$ satisfies
$\rho_{k}^{(j)}\geq\sqrt{\frac{M}{6(1-8\eta)}},$
we have
$f(x_{k}+d_{k}^{(j)})-f(x_{k})$ $\leq\nabla f(x_{k})^{T}d_{k}^{(j)}+\frac{1}{2}\left(d_{k}^{(j)}\right)^{T}\nabla^{2}f(x_{k})d_{k}^{(j)}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
$\leq\nabla f(x_{k})^{T}d_{k}^{E}+\frac{1}{2}\left(d_{k}^{E}\right)^{T}\nabla^{2}f(x_{k})d_{k}^{E}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
$\leq-\frac{1}{8\rho_{k}^{(j)}}\epsilon^{3/2}+\frac{M}{48(\rho_{k}^{(j)})^{3}}\epsilon^{3/2}.$
$\leq-\frac{\eta}{\rho_{k}^{(j)}}\epsilon^{3/2},$ (A.6)
where the second inequality is from the optimality, and the third inequality is because of (2.9b) and (2.9d). As a result, the sufficient descent is satisfied.
When $\|\nabla f(x_{k})\|>\epsilon$, we have three possible outcomes:
  • The first case is
    $\lambda_{\min}(\nabla^{2}f(x_{k}))\leq-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$
    The analysis is almost the same as that in the above case, except that we need to replace $\epsilon$ with $\|\nabla f(x_{k})\|$.
  • The second case is
    $\lambda_{\min}(\nabla^{2}f(x_{k}))\geq\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2},$
    and we need to divide this case into two subcases. The first one is that the dual variable $\lambda_{k}^{(j)}>0$, then it follows $\left\|d_{k}^{(j)}\right\|=\frac{1}{2\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{1/2}$. Moreover, once the parameter $\rho_{k}^{(j)}$ satisfies
    $\rho_{k}^{(j)}\geq\sqrt{\frac{M}{6(1-8\eta)}},$
    we have
    $f(x_{k}+d_{k}^{(j)})-f(x_{k})$ $\leq\nabla f(x_{k})^{T}d_{k}^{(j)}+\frac{1}{2}(d_{k}^{(j)})^{T}\nabla^{2}f(x_{k})^{(j)}d_{k}^{(j)}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
    $=-\frac{1}{2}(d_{k}^{(j)})^{T}\nabla^{2}f(x_{k})^{(j)}d_{k}^{(j)}-\lambda_{k}^{(j)}\left\|d_{k}^{(j)}\right\|^{2}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
    $\leq-\frac{1}{2}\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}\left\|d_{k}^{(j)}\right\|^{2}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
    $\leq-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}.$ (A.7)
    The second subcase is the dual variable $\lambda_{k}^{(j)}=0$. Once the parameter $\rho_{k}^{(j)}$ satisfies
    $\rho_{k}^{(j)}\geq\sqrt{\frac{M}{8\xi}},$
    we have
    $\|\nabla f(x_{k}+d_{k}^{(j)})\|$ $\leq\|\nabla^{2}f(x_{k})^{(j)}d_{k}^{(j)}+\nabla f(x_{k})\|+\frac{M}{2}\left\|d_{k}^{(j)}\right\|^{2}$
    $=\frac{M}{2}\left\|d_{k}^{(j)}\right\|^{2}$
    $=\frac{M}{8(\rho_{k}^{(j)})^{2}}\|\nabla f(x_{k})\|$
    $\leq\xi\|\nabla f(x_{k})\|.$ (A.8)
    It is easy to see the function value is decreasing.
  • The third case is
    $-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$
    Similarly, on one hand if $\lambda_{k}^{(j)}>0$, then $\left\|d_{k}^{(j)}\right\|=\frac{1}{4\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{1/2}$. Once the parameter $\rho_{k}^{(j)}$ satisfies
    $\rho_{k}^{(j)}\geq\sqrt{\frac{M}{12(1-32\eta)}},$
    we have
    $f(x_{k}+d_{k}^{(j)})-f(x_{k})\leq$ $\nabla f(x_{k})^{T}d_{k}^{(j)}+\frac{1}{2}(d_{k}^{(j)})^{T}\nabla^{2}f(x_{k})^{(j)}d_{k}^{(j)}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
    $=$ $-\frac{1}{2}(d_{k}^{(j)})^{T}\nabla^{2}f(x_{k})^{(j)}d_{k}^{(j)}-\lambda_{k}^{(j)}\left\|d_{k}^{(j)}\right\|^{2}$
    $-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}\left\|d_{k}^{(j)}\right\|^{2}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
    $\leq$ $-\frac{1}{2}\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}\left\|d_{k}^{(j)}\right\|^{2}+\frac{M}{6}\left\|d_{k}^{(j)}\right\|^{3}$
    $\leq$ $-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}.$ (A.9)
    On the other hand, if $\lambda_{k}^{(j)}=0$, once the parameter $\rho_{k}^{(j)}$ satisfies
    $\rho_{k}^{(j)}\geq\sqrt{\frac{M}{32\xi-8}},$
    we have
    $\|\nabla f(x_{k}+d_{k}^{(j)})\|$ $\leq\|\nabla^{2}f(x_{k})^{(j)}d_{k}^{(j)}+\nabla f(x_{k})\|+\frac{M}{2}\left\|d_{k}^{(j)}\right\|^{2}$
    $=\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}\left\|d_{k}^{(j)}\right\|+\frac{M}{2}\left\|d_{k}^{(j)}\right\|^{2}$
    $=\frac{1}{4}\|\nabla f(x_{k})\|+\frac{M}{32(\rho_{k}^{(j)})^{2}}\|\nabla f(x_{k})\|$
    $\leq\xi\|\nabla f(x_{k})\|.$ (A.10)
    Also, from the last but one line of (A.9), we have $f(x_{k}+d_{k}^{(j)})-f\left(x_{k}\right)\leq 0$.
So far we have proved (4.1). For (4.2), noting that $\sigma_{k}^{(j)}r_{k}^{(j)}\leq\frac{1}{4}$ in all cases of Strategy 4.1, we have
$\|\nabla f(x_{k}+d_{k}^{(j)})\|$ $\leq\left(\frac{M}{2}(r_{k}^{(j)})^{2}+\sigma_{k}^{(j)}r_{k}^{(j)}\right)\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|$ (A.11)
$\leq\left(\frac{M}{2}(r_{k}^{(j)})^{2}+\frac{1}{4}\right)\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|.$
Thus when $\rho_{k}^{(j)}\to\rho_{\max}$, (4.2) is guaranteed hold.
Now we consider (4.3), and it holds that
$\|\nabla f(x_{k}+d_{k}^{(j)})\|$ $\leq\frac{M}{2}\|d_{k}^{(j)}\|^{2}+\|\left(\sigma_{k}^{(j)}I+\lambda_{k}^{(j)}I\right)d_{k}^{(j)}\|$
$\leq\left(\frac{M(r_{k}^{(j)})^{2}}{2}+1\right)\|\nabla f(x_{k})\|$
$\leq\left(\frac{M}{8(\rho_{k}^{(j)})^{2}}+1\right)\|\nabla f(x_{k})\|.$
Thus when $\rho_{k}^{(j)}\to\rho_{\max}$, (4.3) holds.
In summary, we show in all cases, the inner loop safely terminates as $\rho_{k}^{(j)}$ reaches a bounded constant. ∎
Proof to Lemma 4.2
Proof.
We still prove the result by mathematical induction. The bound obviously holds for $i=0$, and we assume that it also holds for $i=k$. Now let’s check $i=k+1$.
We first check the case $\|\nabla f(x_{k})\|\geq\epsilon$. From (4.2), we have
$\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$ (A.12)
If $\|\nabla f(x_{k})\|\geq\left(\frac{\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\eta}\right)^{2/3}$, from (4.1), we must have $\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|$. Otherwise,
$f(x_{k}+d_{k}^{(j)})\leq f(x_{k})-\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}\leq f(x_{0})-\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}<f^{*},$
and we come to a contradiction. If $\epsilon\leq\|\nabla f(x_{k})\|<\left(\frac{\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\eta}\right)^{2/3}$, we consider two different situations. The first one is $\lambda_{k}^{(j)}<\frac{2M}{3}r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$, then from (A.12) we have
$\|\nabla f(x_{k}+d_{k}^{(j)})\|$ $\leq\left(\xi+\frac{2M}{3}(r_{k}^{(j)})^{2}\right)\|\nabla f(x_{k})\|$
$\leq\left(\xi+\frac{M}{6(\rho_{k}^{(j)})^{2}}\right)\|\nabla f(x_{k})\|$
$\leq\left(\xi+\frac{M}{6\rho_{\min}^{2}}\right)\|\nabla f(x_{k})\|$
$<\left(\xi+\frac{M}{6\rho_{\min}^{2}}\right)\left(\frac{\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\eta}\right)^{2/3}.$
The second one is $\lambda_{k}^{(j)}\geq\frac{2M}{3}r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$, where from (2.11) we have
$f(x_{k}+d_{k}^{(j)})-f(x_{k})\leq-\frac{\lambda_{k}^{(j)}}{4r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}}(r_{k}^{(j)})^{3}\|\nabla f(x_{k})\|^{3/2}.$
Rearrange items we have
$\lambda_{k}^{(j)}r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}\leq\frac{4\left(f(x_{0})-f^{*}\right)}{r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}}.$
Plugging the above inequality into (A.12) yields that
$\|\nabla f(x_{k}+d_{k}^{(j)})\|$ $\leq\xi\|\nabla f(x_{k})\|+\frac{4\left(f(x_{0})-f^{*}\right)}{r_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}}$
$\leq\xi\|\nabla f(x_{k})\|+\frac{16\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\|\nabla f(x_{k})\|^{1/2}}.$
Since $\epsilon\leq\|\nabla f(x_{k})\|<\left(\frac{\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\eta}\right)^{2/3}$, with the same argument as in the fixed case and some basic calculation, we have
$\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\max\left\{\frac{32\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\sqrt{\epsilon}},(\xi+16\eta)\left(\frac{\rho_{\max}\left(f(x_{0})-f^{*}\right)}{\eta}\right)^{2/3}\right\},$
and thus the claim follows.
Till now we have finished the proof for the case $\|\nabla f(x_{k})\|\geq\epsilon$. Regarding the case $\|\nabla f(x_{k})\|<\epsilon$, the only difference in the proof is that (A.12) should be replaced by
$\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\epsilon+\lambda_{k}^{(j)}\frac{1}{2\rho_{k}^{(j)}}\epsilon^{1/2},$
and (2.11) should be replaced by
$f(x_{k}+d_{k}^{(j)})\leq f(x_{k})-\left(\frac{\lambda_{k}^{(j)}}{2r_{k}^{(j)}\epsilon^{1/2}}-\frac{M}{6}\right)(r_{k}^{(j)})^{3}\epsilon^{3/2}.$
The rest discussion on $\lambda_{k}^{(j)}$ is the same as the case of $\|\nabla f(x_{k})\|\geq\epsilon$, and thus the conclusion follows. ∎

### v4

@@
-Beyond Nonconvexity: A Universal Trust-Region Method with New Analyses
+Yuntian Jiang School of Information Management and Engineering
 Abstract
-The trust-region (TR) method is renowned historically for its robustness in nonconvex problems and extraordinary numerical performance, but the study of its performance in convex optimization is somehow limited.
+‘The trust-region (TR) method is renowned historically for its robustness in nonconvex problems and extraordinary numerical performance, but the study of its performance in convex optimization is somehow limited.
 This paper complements the existing literature by presenting a universal trust-region method that simultaneously incorporates the quadratic regularization and ball constraint.
@@
 Recall that the TR method operates on the local quadratic approximation of objective function $f:\mathbb{R}^{n}\to\mathbb{R}$ within a TR ball constraint:
-$\min_{d\in\mathbb{R}^{n}}$ $m_{k}(d):=\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}\nabla^{2}f(x_{k})d,$ (1.1)
+$\min_{d\in\mathbb{R}^{n}}$ $m_{k}(d)=\nabla f(x_{k})^{T}d+\frac{1}{2}d^{T}\nabla^{2}f(x_{k})d,$ (1.1)
 $\operatorname{s.t.}$ $\|d\|\leq\Delta_{k}.$
@@
 $\min_{d\in\mathbb{R}^{n}}$ $\frac{1}{2}d^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\left\|\nabla f(x_{k})\right\|^{1/2}I\right)d+\nabla f(x_{k})^{T}d$ (1.5)
-s.t. $\|d\|\leq\Delta_{k}:=r_{k}\|\nabla f(x_{k})\|^{1/2},$
+s.t. $\|d\|\leq\Delta_{k}=r_{k}\|\nabla f(x_{k})\|^{1/2},$
 where $\sigma_{k},r_{k}>0$ are hyperparameters.
@@
 In this paper, we consider the following unconstrained optimization problem
-$\min_{x\in\mathbb{R}^{n}}\quad f(x),$ (2.1)
+$\min_{x\in\mathbb{R}^{n}}\qquad f(x),$ (2.1)
 where $f:\mathbb{R}^{n}\to\mathbb{R}$ is twice differentiable and is bounded below, i.e., $f^{*}:=\inf_{x\in\mathbb{R}^{n}}f(x)>$-$\infty$.
@@
 The Hessian $\nabla^{2}f(x)$ of the objective function is Lipschitz continuous with constant $M>0$, i.e.,
-$\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq M\|x-y\|\quad\forall x,y\in\mathbb{R}^{n}.$ (2.5)
+$\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq M\|x-y\|\qquad\forall x,y\in\mathbb{R}^{n}.$ (2.5)
 As a consequence, Assumption 2.1 implies the following results.
@@
 Property 2.2 (Function-or-Stationarity-Decrease).
-For some $10<\xi<$ 1, $\kappa>0$, the step $d_{k}$ either decreases the value of the objective function or decreases the gradient norm sufficiently, that is for each iteration $k$,
+For some $0<\xi<1$, $\kappa>0$, the step $d_{k}$ either decreases the value of the objective function or decreases the gradient norm sufficiently, that is for each iteration $k$,
 $f(x_{k}+d_{k})-f(x_{k})\leq-\frac{\kappa}{\sqrt{M}}\|\nabla f(x_{k})\|^{3/2}\ \text{ or }\ \|\nabla f(x_{k}+d_{k})\|\leq\xi\|\nabla f(x_{k})\|.$ (2.8)
@@
 $=-\frac{1}{2}\left(\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\right)\|d_{k}\|^{2}$
-$\qquad-\frac{1}{2}d_{k}^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
+$\hskip 17.00024pt-\frac{1}{2}d_{k}^{T}\left(\nabla^{2}f(x_{k})+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}I+\lambda_{k}I\right)d_{k}+\frac{M}{6}\|d_{k}\|^{3}$
 $\leq-\frac{1}{2}\left(\lambda_{k}+\sigma_{k}\|\nabla f(x_{k})\|^{1/2}\right)\|d_{k}\|^{2}+\frac{M}{6}\|d_{k}\|^{3}$
@@
 Noticing $\lambda_{k}\geq 0$, it is easy to validate that
-$\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\cdot r_{k}^{3}\geq\frac{1}{81\sqrt{M}}\quad\text{and}\quad\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}=\frac{1}{6}.$
+$\left(\frac{1}{2r_{k}}\cdot\frac{\lambda_{k}}{\|\nabla f(x_{k})\|^{1/2}}+\frac{\sigma_{k}}{2r_{k}}-\frac{M}{6}\right)\cdot r_{k}^{3}\geq\frac{1}{81\sqrt{M}}\qquad\text{and}\qquad\frac{M}{2}r_{k}^{2}+\sigma_{k}r_{k}=\frac{1}{6}.$
 Substituting the above inequalities into Lemma 2.3 and Lemma 2.4 completes the proof. ∎
@@
 $\|\nabla f(x_{k}+d_{k})\|\leq\left(1/\xi\right)\cdot\|\nabla f(x_{k})\|,$ (3.6)
-where $10<\xi<$ 1 is the same as that defined in Property 2.2.
+where $0<\xi<1$ is the same as that defined in Property 2.2.
 The above property is a safeguard for the iterates so that the gradient is bounded even in the case where $\lambda_{k}\neq 0$, cf.
@@
 Denoting the sequence generated by the algorithm as $\{x_{k}\}$, we assume that $x_{k}\to x^{*}$, $k\to+\infty$, where $x^{*}$ satisfies
-$\nabla f(x^{*})=0,\quad\nabla^{2}f(x^{*})\succeq\mu I\succ 0.$ (3.11)
+$\nabla f(x^{*})=0,\qquad\nabla^{2}f(x^{*})\succeq\mu I\succ 0.$ (3.11)
 First, we prove that under Assumption 3.2, when $k$ is sufficiently large, the trust-region constraint (2.9a) will be inactive in reminiscences of the classical results.
@@
 Given $\frac{1}{4}<\xi<1$, $0<\eta<\frac{1}{32}$, the step $d_{k}^{(j)}$ satisfies
-{$f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}\text{ or }\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\epsilon^{3/2},\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon,\end{aligned}\right.$ (4.1)
+$\small\left\{\begin{aligned} f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{3/2}\text{ or }\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ f(x_{k}+d_{k}^{(j)})-f(x_{k})&\leq-\frac{\eta}{\rho_{k}^{(j)}}\epsilon^{3/2},\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon,\end{aligned}\right$.
+(4.1)
 where $\eta$ and $\rho_{k}^{(j)}$ are defined in Algorithm 2. If $f$ is nonconvex, the step $d_{k}^{(j)}$ additionally satisfies
-$&\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\epsilon+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon.\end{aligned}\right.$ (4.2)
+$\small\left\{\begin{aligned} &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\|\nabla f(x_{k})\|+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|\geq\epsilon,\\ &\|\nabla f(x_{k}+d_{k}^{(j)})\|\leq\xi\epsilon+\lambda_{k}^{(j)}\|d_{k}^{(j)}\|,\ &\text{if}\ \|\nabla f(x_{k})\|<\epsilon.\end{aligned}\right$.
+(4.2)
 If $f$ is convex, the step $d_{k}^{(j)}$ should satisfy
@@
 $\lambda_{\min}(\nabla^{2}f(x_{k}))\geq\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$
-$-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$ $\sigma_{k}^{(j)}=\rho_{k}^{(j)},\ r_{k}^{(j)}=1/4\rho_{k}^{(j)}$
+$\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}$ $\sigma_{k}^{(j)}=\rho_{k}^{(j)},\ r_{k}^{(j)}=1/4\rho_{k}^{(j)}$
 $\|\nabla f(x_{k})\|<\epsilon$ $\lambda_{\min}(\nabla^{2}f(x_{k}))>-\rho_{k}^{(j)}\epsilon^{1/2}$
@@
 As shown in Lemma 4.1, the iterates produced by Algorithm 2 will end up satisfying Property 4.1, therefore we can classify them into the following two categories
-$\mathcal{F}_{p_{s}}$ =$=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\max\left\{\|\nabla f(x_{k})\|^{3/2},\epsilon^{3/2}\right\}\right\},$ (4.5)
-$\mathcal{G}_{p_{s}}$ =$=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\},$
+$\mathcal{F}_{p_{s}}$ $=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\max\left\{\|\nabla f(x_{k})\|^{3/2},\epsilon^{3/2}\right\}\right\},$ (4.5)
+$\mathcal{G}_{p_{s}}$ $=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\},$
 and denote $x_{p_{s}}$ as the first iteration satisfying (2.2) and (2.3).
@@
 f(xk)Ga,where
-[…]
 […]
@@
 For the case where the objective function is convex, we are satisfied with points such that the norm of the gradient less than $\epsilon$, then the two sets becomes
-$\mathcal{F}_{p_{s}}$ =$=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}\right\},$
-$\mathcal{G}_{p_{s}}$ =$=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\}.$
+$\mathcal{F}_{p_{s}}$ $=\left\{k\leq p_{s}:f(x_{k})-f(x_{k+1})\geq\frac{\eta}{\rho_{\max}}\|\nabla f(x_{k})\|^{3/2}\right\},$
+$\mathcal{G}_{p_{s}}$ $=\left\{k\leq p_{s}:\|\nabla f(x_{k+1})\|\leq\xi\|\nabla f(x_{k})\|\right\}.$
 The following analysis is the same as Algorithm 1 with Strategy 2.1. We directly provide the convergence result here.
@@
 To enable efficient routines for trust-region subproblems, we implement two options.
-The first option utilizes the standard Cholesky factorization [2, Algorithm 4.3] and uses a hybrid bisection and Newton method to find the dual variable [32, 33].
+The first option utilizes the standard Cholesky factorization [28, Algorithm 4.3] and uses a hybrid bisection and Newton method to find the dual variable [34, 35].
 When using this option, we name the method after UTR.
@@
 Motivated from Curtis and Wang [11] and Cartis et al.
-[1, Chapter 10], we use the Lanczos method with inexactness of subproblem solutions.
+[6, Chapter 10], we use the Lanczos method with inexactness of subproblem solutions.
 CUTEst benchmarks
@@
 [4]) and variants of the regularized Newton method [24].
-We implement the regularized Newton method (RegNewton) with fixed regularization $\sigma_{k}\in\{10^{-3},5\times 10^{-4}\}$ and an adaptive version following Mishchenko [22, Algorithm 2.3] that is named after RegNewton-AdaN+.
+We implement the regularized Newton method (RegNewton) with fixed regularization $\sigma_{k}\in\{10^{-3},5\times 10^{-4}\}$ and an adaptive version following Mishchenko [24, Algorithm 2.3] that is named after RegNewton-AdaN+.
 Figure 1: Logistic regression on LIVSVM instances: a4a (left) and w8a (right).
@@
 Matrix Completion
-A final experiment was conducted for matrix completion problems as a comparison to a recent TR method, CAT [20].
-Briefly, the matrix completion task involves reconstructing a partially observed matrix $D\in\mathbb{R}^{n_{1}\times n_{2}}$.
-The goal is to recover the original dataset $D$ by a partially observed set $\Omega=\{(i,j)\}$ and the low-rank representation $D=PQ^{T}$, where $P\in\mathbb{R}^{n_{1}\times r}$ and $Q\in\mathbb{R}^{n_{2}\times r}$ with $r\ll n_{1}$ and $r\ll n_{2}$.
-When predicting the power consumption, $n_{1},n_{2}$ are number of features and data points, respectively.
-Allow a baseline regression model:
+A final experiment is conducted on matrix completion problems motivated by a recent TR method, CAT [19].
+The goal is to recover a power consumption matrix $D\in\mathbb{R}^{n_{1}\times n_{2}}$ using a partially observed set $\Omega=\{(i,j)\}$ and the low-rank representation $D=PQ^{T}$, where $P\in\mathbb{R}^{n_{1}\times r}$ and $Q\in\mathbb{R}^{n_{2}\times r}$ with $r\ll n_{1}$ and $r\ll n_{2}$.
+In the above $n_{1}$ represents the number of measurements taken per day within a 15 mins interval and $n_{2}$ represents the number of days.
+We allow the same baseline estimate as [19]:
 $d_{i,j}=\mu+r_{i}+c_{j},$ (5.2)
-where $\mu$ represents the average of all observed measurements, $r_{i},c_{j}$ captures the deviations from $i$ and $j$. We focus on the following regularized problem:
+where $\mu$ represents the average of all observed measurements, $r_{i},c_{j}$ captures the deviations during time $i$ and day $j$. We focus on the following regularized problem:
 $\min_{r,c,p,q}\sum_{(i,j)\in\Omega}(D_{i,j}-\mu-r_{i}-c_{j}-p_{i}q_{j}^{T})^{2}+\lambda(r_{i}^{2}+c_{j}^{2})+\lambda(\|p_{i}\|^{2}_{2}+\|q_{j}\|^{2}_{2})$
-to fit the observed data under (5.2).
+to fit the observed data under (5.2), where $D_{i,j}$ denotes the observed data recorded at time $i$ and day $j$.
 The dataset used for this study is sourced from Ausgrid [19] that deals with the matrix size of $48\times 30$, $r=9$, all algorithms are terminated at $\|\nabla f(x_{k})\|\leq 10^{-7}$.
@@
 While smaller $\lambda$ induces harder instances, iUTR has the best performance among competing algorithms in all instances.
-Unfortunately, the CAT method only converges to inaccuracy solution with $\|\nabla f(x_{k})\|\leq 10^{-5}$; hence we only include its results when $\lambda=10^{-2}$.
+Since the CAT method only converges to $\epsilon$-FOSPs according to its theory, we only include its results when $\lambda=10^{-2}$.
 Table 4: Performance of different algorithms on Matrix Completion [19].
@@
   • The third case is
-    $-\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$
+    $\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}<\lambda_{\min}(\nabla^{2}f(x_{k}))<\rho_{k}^{(j)}\|\nabla f(x_{k})\|^{1/2}.$
     Similarly, on one hand if $\lambda_{k}^{(j)}>0$, then $\left\|d_{k}^{(j)}\right\|=\frac{1}{4\rho_{k}^{(j)}}\|\nabla f(x_{k})\|^{1/2}$. Once the parameter $\rho_{k}^{(j)}$ satisfies